### v1

# Symplectic torus actions on complete intersections

Abstract
In [20] the author classified complete intersections of dimension $8k$ admitting a Hamiltonian circle action, under a certain assumption on the fixed point set. In this paper, we improve this result in two directions. Firstly, in dimension $8$ we remove the assumption on the fixed point set. Secondly, in any dimension we prove the result under an analogous assumption on the fixed point set. We also give some applications of our methods to the unimodality of Betti numbers of symplectic manifolds having a Hamiltonian $S^{1}$-action, and discuss the relation to symplectic rationality problems.
1 Introduction
To state our results, we set the following notation for complete intersections. Let $X_{m}(d_{1},\ldots,d_{k})$ denote a smooth complete intersection in $\mathbb{CP}^{k+m}$ with multidegree $(d_{1},\ldots,d_{k})$. Here, $m$ is equal to the complex dimension, and $k$ the complex codimension.
Since complete intersections of dimension at least $4$ are simply connected, there is no distinction between Hamiltonian and symplectic circle actions. To shorten the statement of results, we keep the understanding that a complete intersection $(M,\omega)$ is a closed symplectic manifold diffeomorphic to a complete intersection. For some results appearing later in the article, some more restrictions on the symplectic form are needed and these are stated explicitly.
The major motivating result of this paper is a theorem of Benoist which gave a complete characterization of smooth complete intersections whose holomorphic automorphism groups are infinite [1, Theorem 3.1]. The aim of this work is to obtain a symplectic analogue of this theorem. A comprehensive smooth analogue in dimension $6$ was given by Dessai and Wiemeler [8].
The first main result is to remove the technical assumption on the fixed point set in our previous result on complete intersections [20, Theorem 1.1] in dimension 8.
Theorem 1.1.
Suppose that $(M,\omega)$ is an $8$-dimensional complete intersection with a symplectic circle action. Then, $M$ is diffeomorphic to either $X_{4}(1),X_{4}(2)$ or $X_{4}(2,2)$.
The main issue to overcome in the proof is to deal with $4$-dimensional fixed components. For this we apply a result of $4$-dimensional symplectic geometry due to Ohta and Ono (building on fundamental work of Gromov, McDuff, and Taubes), to deduce information about their intersection form, and deduce via localisation formulas, information about the intersection form of $M$.
In the next part of the paper, we prove a classification result for symplectic circle actions on complete intersections under a condition on the fixed point set analogous to [20, Theorem 1.1], therefore completely removing the assumption on the dimension.
Theorem 1.2.
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action with no fixed components $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$.
Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
Theorem 1.2 has the following corollary.
Corollary 1.3.
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action having only isolated fixed points. Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
Projective spaces and quadrics of any dimension admit a Hamiltonian torus action with isolated fixed points. The underlying smooth $4$-manifolds of the complete intersections $X_{2}(2,2),$ or $X_{2}(3),$ admit a symplectic form with a toric action, since both are diffeomorphic to the blow-up of $\mathbb{CP}^{2}$ in a sequence of points.
Hence, the only unknown case remaining from Theorem 1.3 is $X_{2m}(2,2)$ for $m\geq 2$. We are able to deal with this when the action is GKM, and the symplectic form is homotopic to the Fubini-Study form, which we denote $\omega_{FS}$. Thus obtaining in Theorem 4.2, a Benoist type “if and only if” classification of complete intersections having such symplectic torus actions, precisely when the holomorphic automorphism group, denoted $Aut(M)$ is infinite. The proof uses results from [14], to evaluate the integral of $c_{1}$ along a hypothetical toric $1$-skeleton Poincaré dual to $c_{n-1}$, and thus obtaining a contradiction with the known Chern numbers of the intersection of two quadrics.
Theorem 1.4.
Let $M$ be a complete intersection. Then $(M,\omega_{FS})$ has a GKM symplectic torus action $\iff\;\;|Aut(M)|=\infty$.
In the final section of the paper, we prove some new results regarding the unimodality of Betti numbers of a closed symplectic manifold admitting a Hamiltonian circle action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian circle action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [15, Problem 4.3]. For Kähler manifolds the property follows from the hard Lefschetz theorem. [15, Problem 4.3] also asked about the sequence of odd Betti numbers, but the results of this paper are only relevant for the even case.
In [5] Cho and Kim gave an affirmative answer to [15, Problem 4.3 (1)] when $M$ is $8$-dimensional and the action has isolated fixed points. In Theorem 5.1 we prove the unimodality of even Betti numbers for $8$-dimensional manifolds having a Hamiltonian $S^{1}$-action with no extremal, 4-dimensional fixed components.
Theorem 1.5.
Suppose that $M$ is a closed symplectic $8$-manifold having a Hamiltonian $S^{1}$-action, and neither $M_{\min}$ or $M_{\max}$ are $4$-dimensional. Then $b_{2}(M)\leq b_{4}(M)$.
In Theorem 5.1 we prove that for a closed symplectic $12$-manifold with $b_{2}(M)=1$, satisfying a similar assumption on the fixed point set the even Betti numbers are unimodal. To the author‘s knowledge this is new under the assumption of isolated fixed points.
Furthermore, we give versions of the above results in general dimensions, generalizing an inequality of Cho [3] on the Betti numbers of a manifold admitting a Hamiltonian $S^{1}$-action with isolated fixed points , to more general fixed point sets. Finally, we combine this result with the Rokhlin-Ochanine theorem to obtain a final application.
Acknowledgments. I would like to thank Leonor Godinho and Silvia Sabatini for helpful discussions, in particular explaining to me the results of [14], which allowed me to deal with the intersection of two quadrics in the GKM case. I presented some of the results about unimodality of Betti numbers in the workshop “Perspectives in equivariant topology” in Cologne September 2022, I would like the thank those present for interesting discussions, in particular Liat Kessler and Susan Tolman. I would like to thank Dmitri Panov for collaborations and discussions which influenced the overall approach of the project, and for comments on a draft version.
The author was supported by SFB-TRR 191 grant Symplectic Structures in Geometry, Algebra and Dynamics funded by the Deutsche Forschungsgemeinschaft.
2 Preliminaries
2.1 Hamiltonian circle actions
Firstly, we recall the definition of Hamiltonian torus action on a symplectic manifold. Let $\mathfrak{t}$ denote the Lie algebra of the torus $T^{k}=T$, and $\mathfrak{t}^{*}$ its dual. Recall that if a manifold $M$ has a smooth $T$-action, then for every $t\in\mathfrak{t}$ we can naturally associate an invariant vector field on $M$; $X_{t}$, tangent to the orbits of the action of the one-parameter family associated to $t$.
Definition 2.1.
Let $(M,\omega)$ be a closed symplectic manifold, with a smooth $T^{k}$-action by symplectomorphims. Then, the action is called Hamiltonian if there is a smooth function $\mu:M\rightarrow\mathfrak{t}^{*}$ such that
$\omega(X_{t},\cdot)=-d(\mu(t)),$
for all $t\in\mathfrak{t}$.
$\mu$ is called the moment map of the action. The fixed point set of the action, denoted $M^{T}$ is equal to the critical set of $\mu$, and is a collection of symplectic submanifolds.
When $k=1$ the moment map is ussually referred to as the Hamiltonian and denoted $H:M\rightarrow\mathbb{R}$. The Hamiltonian $H$ is a perfect Morse-Bott function [17], and a perfect Morse function if and only if $M^{S^{1}}$ is finite.
Let $G\subset S^{1}$ be a subgroup. The subset $M^{G}=\{x\in M|\;g.p=p\;\forall g\in G\}$ is an invariant symplectic submanifold for which the action is Hamiltonian, with Hamiltonian $H|_{M^{G}}$. This will be used in the proof of Theorem 1.1. Connected components of $M^{G}$ are called isotropy submanifolds.
For a fixed component $F$ denote by $\lambda_{F}$ half the Morse-Bott index of $H$ along $F$. It holds that $\lambda_{F}$ is also equal to the the number of strictly negative weights of the $S^{1}$-action along $F$ counted with multiplicity. We will need the following localisation formula for the Betti numbers.
Theorem 2.2.
[17] Let $(M,\omega)$ be a closed symplectic manifold having a Hamiltonian circle action. Then for each $i$,
$b_{i}(M)=\sum_{F\subset M^{S^{1}}}b_{i-2\lambda_{F}(M)}(F),$
where the sum runs over connected components of the fixed point set.
The following localisation formula is a consequence of the equivariant version of the Atiyah-Singer index theorem. The formula we will use is derived in [12, Section 5.8]. We use $\sigma(M)$ to denote the signature of $M$, which is equal to zero for manifolds of dimension not divisible by $4$ and equal to $1$ for a single point.
Theorem 2.3.
Let $(M,\omega)$ be a closed symplectic manifold having a Hamiltonian $S^{1}$-action. Then
$\sigma(M)=\sum_{F\subset M^{S^{1}}}(-1)^{\lambda_{F}}\sigma(F).$
Proof.
This formula follows from the formula of [12, Section 5.8], expressing the signature as a sum over the fixed point components.
We will need the following result which was proved in [20].
Theorem 2.4.
[20, Proposition 3.1, Lemma 3.2] Suppose that $M$ is a closed symplectic manifold of dimension $8k$. Suppose that $M$ has a Hamiltonian $S^{1}$-action such that each component of the fixed point set is an isolated fixed point or has dimension $2\mod 4$. Suppose also that $b_{i}(M)=b_{i}(\mathbb{CP}^{4k})$ for $i\neq 4k$. Then the intersection form on $H^{4k}(M,\mathbb{Z})$ is positive definite.
2.2 Applications of Duistermaat-Heckman
We begin by recalling the Dusitermaat-Heckman theorem.
Theorem 2.5.
Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action, with Hamiltonian $H$. Let $c$ be a regular value of $H$. Then
$\frac{d}{dt}[\omega_{t}]=e(H^{-1}(M)).$
Here $H^{-1}(M)$ is considered as an (orbifold) principal $S^{1}$-bundle over the reduced symplectic orbifold $M_{t}$. Finally, $e$ denotes the Euler class of this bundle in $H^{2}(M_{t},\mathbb{R})$.
Here we denote by $M_{\min}$ and $M_{\max}$ the fixed point components on which the Hamiltonian obtains its minimum and maximum respectively. Recall that they are both connected, symplectic submanifolds.
We call a fixed point component $Y$ internal if $H(M_{\min})<H(Y)<H(M_{\max})$.
Lemma 2.6.
Suppose that $M$ is a closed symplectic manifold with a Hamiltonian circle action. Suppose that $M_{\min}$ consists of an isolated fixed. Suppose that $Y$ is an internal fixed point component such that $H(Y^{\prime})\geq H(Y)$ for any other internal fixed component $Y^{\prime}$. Then, $\lambda_{Y}=1$.
Proof.
This is an immediate consequence of [29, Lemma 3.1]. ∎
We will need the following result which follows from results of Ono.
Lemma 2.7.
[14, Lemma 5.2] Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action and $b_{2}(M)=1$. Then after possibly rescaling the symplectic form by a positive constant we may assume that $c_{1}(M)=[\omega]$.
Proof.
Since $b_{2}(M)=1$, $c_{1}(M)=k[\omega]$ for some $k\in\mathbb{R}$. The result then follows from [25, Theorem 1,Theorem 2].
For Hamiltonian $S^{1}$-actions on Fano varieties, the following normalization for the Hamiltonian goes back to Futaki [11, Theorem 1]. In the case of monotone symplectic manifolds proofs may be found for the semi-free case in [6] and in the general case [14]. In [21, Proposition 1.9] the formula was proven for closed symplectic manifolds with Hamiltonian $S^{1}$-actions, such that $c_{1}$ and $[\omega]$ have the same evaluation on classes in the image of the Hurewicz map $\pi_{2}(M)\rightarrow H_{2}(M)$, this will not be needed in this paper.
Theorem 2.8.
Suppose that $(M,\omega)$ is a closed symplectic manifold with $c_{1}(M)=[\omega]$. Then, after possibly adding a constant to the Hamiltonian, for each fixed component $F$,
$H(F)=-\sum w_{i},$
where $w_{i}$ are the weights along $F$.
There has been a lot of recent progress towards classifying monotone symplectic manifolds with Hamiltonian symmetries, see [4, 2, 14, 21, 28] and the reference contained therein.
The following theorem is due to Ohta and Ona, and builds on fundamental work of Gromov, McDuff and Taubes.
Theorem 2.9.
[26] Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $[\omega]=kc_{1}(N)$ for some $k>0$. Then, $N$ is diffeomorphic to a del Pezzo surface.
The theorem has the following immediate corollary.
Corollary 2.10.
Let $(N,\omega)$ be a closed symplectic $4$-manifold such that $[\omega]=kc_{1}(N)$ for some $k>0$. Then, $b_{2}^{+}(N)=1$.
Proof.
By Theorem 2.9, $N$ is diffeomorphic to a smooth del Pezzo surface. By the classification of del Pezzo surfaces, $N$ is either diffeomorphic to $\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ or the blow up of $\mathbb{CP}^{2}$ in up to $9$ points. One may verify by direct calculation that each of these manifolds satisfies $b_{2}^{+}=1$. ∎
2.3 Complete Intersections
A smooth complete intersection $M=X^{n}\subset\mathbb{CP}^{n+k}$ is a subvariety of projective space cut out by a collection of smooth hypersurfaces in general position. By applying the Lefschetz hyperplane consecutively on the image of the variety via appropriate Veronese embeddings, it follows that complete intersections satisfy $b_{i}(X)=\mathbb{CP}^{n}$ for $i\neq n$. The topological consequences of this fact have a different flavour when $n$ is odd or even. In both cases, there is an important characterization of when a complete intersection is“close to projective space”, which is due to Ewing and Moolgavkar when $n$ is odd and Libgober and Wood when $n$ is even.
In the case when $n$ is odd, the middle degree homology $H^{n}(M,\mathbb{R})$ has the structure of a symplectic vector space given by the intersection product and does not contain cup products of cohomology classes of other degrees. Thus the “odd part” of the cohomology ring of $M$ is essentially determined by $b_{n}(M)\in 2\mathbb{Z}$. A complete characterization of when this part of the cohomology ring vanishes, as it does for projective space, was given by Ewing and Moolgavkar [9]. Let $Q=X_{n}(2)$ denote the quadric $n$-fold.
Theorem 2.11.
[9, Corollary 5] Suppose that $X$ is a complete intersection of odd complex dimension $n$, then if $\chi_{top}(M)=n+1$ then $X\cong\mathbb{CP}^{n}$ or $X\cong Q$.
In the case when $n$ is even, then the intersection form on $H^{n}(M,\mathbb{R})$ is symmetric, and $H^{n}(M,\mathbb{R})$ plays a more interesting role in determining the cohomology ring of $M$. The most fundamental invariant we can extract from the intersection form is the signature $\sigma(M)$. A property satisfied by projective space is that the signature is a equal to the alternating sum of even Betti numbers, as was shown by Libgober and Wood, this almost gives a characterization of projective space amongst complete intersections.
Theorem 2.12.
[19] Let $X$ be a complete intersection with $n$ even. Then
$\sigma(X)=\sum_{k\in\mathbb{Z}}(b_{4k}(X)-b_{4k+2}(X))$
if and only if $X$ is equal to $X_{n}(1),X_{n}(2),X_{2}(3)$ or $X_{n}(2,2)$.
Combining the condition $\sigma(X)=\sum_{k\in\mathbb{Z}}(b_{4k}(X)-b_{4k+2}(X))$ with the Lefschetz hyperplane theorem shows that this is equivalent to the intersection form being positve definite when $4|n$ and $b_{n}^{+}(X)=1$ otherwise.
A beautiful fact, proved by Deligne [7, Page 339] is that when $4|n$ the intersection form of $X_{n}(2,2)$ is the Korkine-Zolotareff form $\Gamma_{n+4}$. That is the lattice of $\mathbb{R}^{n+4}$ generated by $e_{i}+e_{j}$ and $\frac{1}{2}(e_{1}+\ldots+e_{n+4}).$ In particular for $n=4$ the intersection form is isomorphic to the $E_{8}$ lattice. Deligne gives a Dynkin diagram description of the intersection form for $n=2\mod 4$, and it is clear from this description that $b^{+}_{n}(X_{n}(2,2))=1$.
3 Proof of Theorem 1.1
We now proceed to the proof of our main result Theorem 3.4. Before giving the details of the proof, we describe the general approach. It is necessary to make the following definition. We follow the convention that for manifolds with real dimension not divisible by $4$ the signature is 00, and a point has signature $1$.
Definition 3.1.
Suppose that $X$ is a closed manifold. Define
$I_{JR}(X)=\sigma(X)-\sum_{i\in\mathbb{Z}}(b_{4i}(X)-b_{4i+2}(X)),$
we refer to $I_{JR}(X)=0$ to be the JR equation.
The idea of the proof of the following proposition was motivated by a theorem of Jones and Rawnsley [16, Theorem 1.1], which states that if a closed symplectic $(M,\omega)$ has a Hamiltonian circle action with isolated fixed points then it satisfies the JR equation.
In the following Proposition 3.2 we show that the invariant $I_{JR}$ localizes in a simple way over the fixed point set. A form of Proposition 3.2 with additional conditions has been proven by Faber [10, Corollary 6]. Fabers result states that for a manifold with a Hamiltonian circle action such that $I_{JR}(F)=0$, $\forall F\subset M^{S^{1}}$ and all the odd Betti numbers of each $F$ vanish, ones has $I_{JR}(M)=0$. Proposition 3.2 improves on [10, Corollary 6] in three ways. Firstly, it removes the assumption on the odd Betti numbers. Secondly, it shows that only fixed submanifolds of positive dimension divisible by $4$ are relevant. Lastly, the equation of Proposition 3.2 can be useful for an inequality analysis when both sides are not necessarily zero, this is key for the proof of Theorem 5.1.
Proposition 3.2.
Suppose that $(M,\omega)$ is a closed symplectic manifold admitting a Hamiltonian circle action. Then,
$I_{JR}(M)=\sum_{\{F\subset M^{S^{1}}|\dim(F)=4k>0\}}(-1)^{\lambda_{F}}I_{JR}(F).$
Proof.
By Theorem 2.2 and Theorem 2.3 the alternating sum of even Betti numbers and the signature may be be expressed as sums over the fixed points set, in a similar way to the proof of [20, Proposition 3.1]. By Theorem 2.2:
$\sum_{j\in\mathbb{Z}}b_{4j}(M)-\sum_{j\in\mathbb{Z}}b_{4j+2}(M)=\sum_{F\subset M^{S^{1}}}(\sum_{j\in\mathbb{Z}}b_{4j-2\lambda_{F}}(F)-\sum_{j\in\mathbb{Z}}b_{4j+2-2\lambda_{F}}(F)).$
For each component $F$, considering how shifting index changes the sign of an alternating sum gives the equation:
$(\sum_{j\in\mathbb{Z}}b_{4j-2\lambda_{F}}(F)-\sum_{j\in\mathbb{Z}}b_{4j+2-2\lambda_{F}}(F))=(-1)^{\lambda_{F}}(\sum_{j\in\mathbb{Z}}b_{4j}(F)-\sum_{j\in\mathbb{Z}}b_{4j+2}(F)).$
On the other hand:
$\sigma(M)=\sum_{F\subset M^{S^{1}}}(-1)^{\lambda_{F}}\sigma(F).$
Using these formulas to express $I_{JR}(M)$ as a sum over the fixed point set, gives the equation.
$I_{JR}(M)=\sum_{F\subset M^{S^{1}}}(-1)^{\lambda_{F}}I_{JR}(F).$
It remains to show that when $F$ is an isolated fixed point or that $\dim(F)=2\mod 4$, that $I_{JR}(F)=0$. For $F$ an isolated fixed point, the statement is straightforward to prove, recalling that a point has signature $1$.
Suppose $N$ has dimension $2\mod 4$, it follows that $\sigma(N)=0$. So to prove $I_{JR}(N)=0$ it is sufficient to prove
$\sum_{i\in\mathbb{Z}}(b_{4i}(N)-b_{4i+2}(N))=0.$
This follows from Poincarè duality and the assumption that $\dim(N)=2\mod 4$.
The following simple lemma simplifies $I_{JR}$ for $4$-dimensional components.
Lemma 3.3.
Suppose that $N$ is a closed manifold of dimension $4$. Then
$I_{JR}(N)=2b_{2}^{+}(N)-2.$
Proof.
Recall that $b_{2}(N)=b_{2}^{+}(N)+b_{2}^{-}(N)$ , $\sigma(N)=b_{2}^{+}(N)-b_{2}^{-}(N)$ and that $b_{0}(N)=b_{4}(N)=1$. So it follows that
$I_{JR}(N)=(b_{2}^{+}(N)-b_{2}^{-}(N))-2+(b_{2}^{+}(N)+b_{2}^{-})=2b_{2}^{+}(N)-2.$
3.1 Proving the JR equation for $4$-dimensional fixed components
In the below theorem, we complete the main step to classifying $8$-dimensional complete intersections with a symplectic circle action. The most important result which motivates Theorem 3.4 is a result of Tolman [29], which classified Hamiltonian $S^{1}$-actions on $6$-manifolds with $b_{2}=1$. In [13], building on results of [14], substantial progress towards the problem of classifying $8$-dimensional symplectic manifolds having a Hamiltonian torus action with isolated fixed points and $b_{2}=1$ was completed. A sharp classification result into four explicit cases was obtained when the torus is two dimensional and $b_{4}=2$ or equivalently the action has exactly $6$ fixed points [13, Main Theorem]. All of the examples subcircle actions with various fixed components of positive dimension occurring.
Theorem 3.4.
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold admitting a Hamiltonian circle action, and that $b_{2}(M)=1$. Then, the intersection form of $M$ is positive definite.
Proof.
Firstly we note that the conclusion of the theorem is equivalent to showing that $I_{JR}(M)=0$. By Theorem 3.2 and Lemma 3.3 it is sufficient to prove for each $4$-dimensional fixed point component $N$, $b_{2}^{+}(N)=1$.
Let $N$ be a $4$-dimensional fixed component, since the downward flowing normal bundle of $N$ has rank at most 2 (over $\mathbb{C}$), it follows that $0\leq\lambda_{N}\leq 2$. We claim that if $\lambda_{N}$ is 00 or $2$, then by Theorem 2.2 $b_{2}(N)=1$. Since $N$ is a symplectic submanifold $b_{2}(N)>0$. In the localisation sum for the Betti numbers Theorem 2.2, $b_{2}(N)$ either contributes to $b_{2}(M)$ or $b_{6}(M)$ (when $\lambda_{N}=0$ or $2$ respectively) via the localisation formula, but by the assumption of Theorem 3.4 and Poincaré duality $b_{2}(M)=b_{6}(M)=1$. It follows that $b_{2}(N)=1$ in each case. Finally, since $b_{2}(N)=1$ and $N$ is a symplectic submanifold $\int_{N}[\omega_{N}]^{2}>0$ implying that $b_{2}^{+}(N)=1$, proving Theorem 3.4 in this case.
Hence we may assume that $\lambda_{N}=1$, this case is more involved. Firstly this assumption implies by Theorem 2.2 that $M_{\min}$ and $M_{\max}$ are isolated fixed points.
By Theorem 2.2 every other fixed component is an isolated fixed point $p$ with $\lambda_{p}=2$. We claim that such fixed points dont occur. If $H(p)\leq H(N)$ then this follows from Lemma 2.6. If $H(p)\geq H(N)$ then this follows from Lemma 2.6, the case $H(p)\leq H(N)$ is proved similarly.
Hence, there are three fixed components $M_{\min}$, $M_{\max}$ and $N$. By Lemma 2.7 we may assume that $c_{1}(M)=[\omega]$. Since $\dim(N)=4$ and $\lambda_{N}=1$ the normal bundle of $N$ splits as a sum of two line bundles, one upward flowing with respect to the Hamiltonian, and one downward flowing. We denote the splitting by $L_{1}\oplus L_{2}$. Let $w_{1}$ and $w_{2}$ denote the weights of the circle action along $L_{1}$ and $L_{2}$ respectively.
If either of $w_{1}$ or $w_{2}$ has modulus greater than $1$, there is an isotropy $6$-manifold $Q$ such that $Q^{S^{1}}$ has two fixed components. Suppose without loss of generality that $w_{1}<-1$, so that $Q_{\max}=N$, $Q_{\min}=M_{\min}$. Then applying Theorem 2.2 to the Hamiltonian $S^{1}$-action on $Q$ gives $b_{2}(Q)=b_{0}(N)=1$. Then, by Poincaré duality $b_{4}(Q)=1$, then applying Theorem 2.2 again $b_{4}(Q)=b_{2}(N)=1$. Then since the symplectic form of $N$ squares positively we have that $b_{2}^{+}(N)=1$ as required.
Hence, we may assume that $w_{1}=-1$ and $w_{2}=1$. In particular, $H(N)=0$ by Theorem 2.8. To ease notation we set $H(M_{\max})=H_{max}$ and $H_{\min}=H(M_{\min}).$ Then since $M_{\min}$ and $M_{\max}$ are isolated points, all of the four weights along them are non-zero and of the same sign. This implies by Theorem 2.8 that $H(N)-H_{\min}=-H_{\min}\leq-4$ and $H_{\max}-H(N)=H_{\max}\geq 4$.
Considering how the tangent bundle of $M$ splits over $N$ shows that
$[\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+c_{1}(L_{1})+c_{1}(L_{2}).$ (1)
Let $\bar{N}$ be the union of gradient sphere with maximum in $N$. We consider the cycle $N_{c}\subset M_{c}$ which is obtained as the projection of $\bar{N}$ by the quotient map $N_{c}$. Considering levels of the Hamiltonian slightly below $0=H(N)$, and applying the equivariant symplectic neighborhood theorem gives that:
$e(H^{-1}(c))|_{N_{c}}=c_{1}(N_{1})|_{N_{c}}\in H^{2}(N_{c},\mathbb{Z}).$
Note that $N_{H_{\min}}$ is a point since $M_{min}$ is isolated, by the equivariant Darboux threorem, the limit of the class of the reduced symplectic form is zero as the Hamiltonian approaches $H_{\min}$. Applying Duistermaat-Heckman Theorem 2.5 and the above observations yields the equation:
$[\omega]|_{N}=(H(N)-H_{\min})c_{1}(L_{1})=|H_{\min}|c_{1}(L_{1})$
Applying similar consinderations to the upward flowing manifold of $N$ gives the following equation:
$[\omega]|_{N}=(H_{\max}-H(N))c_{1}(L_{2})=|H_{\max}|c_{1}(L_{2})$
Substituting these equations back into equation (1) gives
$[\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+(\frac{1}{|H_{\max}|}+\frac{1}{|H_{\min}|})[\omega]|_{N},$
By, Corollary 2.10 $b_{2}^{+}(N)=1$, proving the theorem.
The following simple example may help the reader to get an overview on the last part of the proof.
Example 3.5.
Consider $\mathbb{CP}^{4}$ with the Hamiltonian $S^{1}$-action.
$z.[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]=[x_{0}:zx_{1}:zx_{2}:zx_{3}:z^{2}x_{4}].$
Then $M_{\min}$ and $M_{\max}$ are isolated and are mapped to $-5$ and $5$ respectively by the moment map of Theorem 2.8. The only other fixed component is $N=\mathbb{CP}^{2}$ with $H(N)=0$ and normal bundle $\mathcal{O}(1)\oplus\mathcal{O}(1)$. Therefore $c_{1}(M)|N=c_{1}(T\mathbb{CP}^{2})+2c_{1}(\mathcal{O}(1))=5H$, showing that $c_{1}(M)|N=|H_{\max}|c_{1}(\mathcal{O}(1))=|H_{\min}|c_{1}(\mathcal{O}(1))$.
We also present the following corollaries which may be of independent interest.
Corollary 3.6.
Suppose that $M$ is an $8$-dimensional closed symplectic manifold with a Hamiltonian circle action and with $b_{2}(M)=1$. Then,
$\int_{M}c_{2}(M)^{2}\geq 0.$
4 Results in arbitrary dimensions
4.0.1 Symplectic circle actions
In the following theorem, we give a classification of the multidegrees of complete intersections admitting a Hamiltonian $S^{1}$-action whose fixed point sets satisfy an analogous assumption to [20, Theorem 1.1].
Theorem 4.1.
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action with no fixed components $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$.
Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
Proof.
Firstly, suppose that $\dim(M)=2n$, with $n$ odd. By applying the Lefschetz hyperplane theorem and Poincaré duality to $M$, all of the even Betti numbers of $M$ are equal to $1$ and all of the odd Betti numbers are 00, except possibly $b_{n}(M)$.
Let $F$ be a fixed submanifold, then by Theorem 2.2, $b_{i}(F)\neq 0$ for at most one odd $i$. Since, by assumption $\dim(F)$ is divisible by $4$ this implies by applying Poincaré duality to $F$ that all of the odd Betti numbers of $F$ vanish, because every odd Betti number appears in at least two degrees. Therefore by Theorem 2.2 the odd Betti numbers of $M$ vanish. Since all of the even Betti numbers are equal to one, it follows that $\chi(M)=n+1$. Hence, by Theorem 2.11 the multidegree of $M$ is (1) or (2).
In the case that $\dim(M)$ is divisible by $4$, then by our assumptions all terms on the right hand side of the equation given in Theorem 3.2 vanish. It follows that
$\sigma(M)=\sum_{k\in\mathbb{Z}}(b_{4k}(M)-b_{4k+2}(M)).$
By Theorem 2.12 the multidegree of $M$ is $(1),(2)$ or $(2,2)$, or $M$ is a cubic surface. ∎
4.0.2 GKM actions
In this section, we strengthen the above results under the additional assumption that the action is GKM and preserves a symplectic form homotopic to the Fubini-Study form. We are able to give a complete characterization as in the classification of infinite automorphism groups due to Benoist [1, Theorem 3.1].
Theorem 4.2.
Let $X$ be a smooth complete intersection of complex dimension greater than $2$, and let $\omega_{0}$ be the restriction of the Fubini-Study form to $X$. Then $X$ admits a GKM torus action preserving a symplectic form that is homotopic to $\omega_{0}$ $\iff$ $|Aut(X)|=\infty$.
Proof.
By Theorem 1.2 and the theorem of Benoist [1, Theorem 3.1], it is sufficient to exclude the existence of such actions on a complete intersection of the form $X_{2m}(2,2)$. Suppose that there is such a GKM action. Set $n=2m$ to be the complex dimension of $X_{2m}(2,2)$. Since $X_{2m}(2,2))$ is Fano it is possible to rescale the hypothetical invariant symplectic form so that $c_{1}=[\omega]$.
A calculation with the adjunction formula and Bezout’s theorem gives that the the Chern number $c_{1}c_{n-1}$ of $X_{n}(2,2)$ is given by
$\frac{n}{2}(n+2)(n-1).$
Moreover, the topological Euler characteristic is
$2(n+2).$
It follows that the the number of edges of the GKM graph is
$n(n+2).$
The collection of spheres corresponding to the edges is Poincaré dual to $c_{n-1}$ [14] and by the adjunction formula $c_{1}$ is $(n-1)$ times the primitive generator of $H^{2}$. Therefore, using the condition $c_{1}=[\omega]$ the integral of $c_{1}$ on each of the spheres is a positive integer multiple of $(n-1)$. Therefore implying that the Chern number $c_{1}c_{n-1}$ is divisible by $n(n+2)(n-1)$, which is a contradiction.
We note finally, that it is possible to obtain a statement covering complex dimension $2$, at the price of requiring that the symplectic form preserved by the GKM action is monotone. Since both $X_{2}(2,2)$ and $X_{2}(3)$ admit Kähler forms preserved by a toric action, this is the strongest possible statement which can cover complex dimension $2$. Since the proof is identical to the proof of Theorem 4.2, it is omitted.
Theorem 4.3.
Let $X$ be a smooth complete intersection and let $\omega_{FS}$ be the restriction of the Fubini-Study form to $X$. Then $X$ admits a GKM torus action preserving a monotone symplectic form that is homotopic to $\omega_{FS}$ $\iff$ $|Aut(X)|=\infty$.
Another, perhaps more elegant but weaker possibility, is to require that the symplectic structure is symplectomorphic to one given by the Fubini-Study symplectic form. This was stated in Theorem 1.4 of the introduction.
5 Unimodality of Betti numbers
Given the generality of Proposition 3.2, it is natural ask whether $I_{JR}(M)=0$ for all closed symplectic manifolds admitting a Hamiltonian circle action. The following example shows that this is not the case.
Consider the standard Hamiltonian $S^{1}$-action on $\mathbb{CP}^{4}$ fixing a point and a hyperplane $H$. Let $S\subset H$ be a smooth degree $4$ hypersurface. Recall that $S$ is a K3 surface with signature $16$ and $b_{2}(S)=22$. Let $M$ be the $S^{1}$-equivariant blow up of $\mathbb{CP}^{4}$ along $S$. Then by Proposition 3.2
$I_{JR}(M)=-I_{JR}(N)=-36.$
In this section, we will present some applications regarding the unimodality of Betti numbers of manifolds admitting a Hamiltonian circle action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian circle action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [15, Problem 4.3]. For Kähler manifolds the property holds from the hard Lefschetz theorem.
In [5] Cho and Kim proved that $8$-dimensional closed symplectic manifolds having a Hamiltonian $S^{1}$-action with isolated fixed points, the even Betti numbers are unimodal. Here we prove the same result under the weaker assumption that neither of the extremal submanifolds are $4$-dimensional.
Theorem 5.1.
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold admitting a Hamiltonian circle action such that, $M_{\min},M_{\max}$ are not $4$-dimensional. Then the even Betti numbers of $M$ are unimodal.
Proof.
By Theorem 3.2 $M$ satisfies:
$I_{JR}(M)=\sum_{\{F\subset M^{S^{1}}|\dim(F)=4k>0\}}(-1)^{\lambda_{F}}I_{JR}(F).$
Since by assumption $M_{\min}$ and $M_{\max}$ are not $4$-dimensional, every $4$-dimensional fixed component $F$, satisfies $\lambda_{F}=1$.
Therefore the above equation becomes:
$I_{JR}(M)=\sum_{\{F\subset M^{S^{1}}|\dim(F)=4\}}-I_{JR}(F).$
By Lemma 3.3 $I_{JR}(F)=2b^{+}(F)-2$, since $F$ is a symplectic $4$-manifold, this gives $I_{JR}(F)>0$. Therefore giving $I_{JR}(M)<0$.
Or, the inequality may me restated as
$\sigma(M)-2+2b_{2}(M)-b_{4}(M)<0.$
Then, since $M$ is symplectic $-b_{4}(M)<\sigma(M)$, substituting this and rearranging gives $1+b_{2}(M)<b_{4}(M)$. Then, since both $b_{2}(M),b_{4}(M)$ are integers this gives the desired inequality.
Motivated by Theorem 5.1, we state a conjecture. Following [18] a closed symplectic manifold $(M,\omega)$ is called symplectically rational if there is a connected symplectic manifold with a semi-free Hamiltonian $S^{1}$-action such that the reduced spaces at two regular levels are symplectomorphic to $(M,\omega)$ and $(\mathbb{CP}^{n},\omega_{FS})$ respectively.
Conjecture 5.2.
Suppose that $M$ is a symplectically rational $8$-manifold admitting a Hamiltonian $S^{1}$-action. Then, the even Betti numbers of $M$ are unimodal.
If a complex rational $4$-fold $M$ has such an algebraic torus action having an algebraic surface $N$ as an extremal component, then $M$ is birational to a weighted projective space bundle over $N$ via a standard GIT argument. Therefore $N$ is stably rational and hence rational [30]. Therefore $N$ contributes 00 to the equation of Proposition 3.2 and the proof of Theorem 5.1 goes through.
We also obtain a similar result for closed $12$-manifolds with $b_{2}=1$. To the authors knowledge this result is new, under the weaker assumption of isolated fixed points.
Proposition 5.3.
Suppose that $(M,\omega)$ is a closed symplectic $12$-manifold admitting a Hamiltonian circle action with no $4$ or $8$ dimensional fixed components, and $b_{2}(M)=1$. Then the even Betti numbers of $M$ are unimodal.
Proof.
By Theorem 3.2 $M$ satisfies the JR equation. Since the modulus of the signature of a symmetric bilinear form is bounded above by the rank of the middle homology group, we have that
$\sigma(M)\leq b_{6}(M).$
Then substituting the JR equation for $\sigma(M)$ gives
$2b_{4}(M)-b_{6}(M)\leq b_{6}(M).$
Which implies that $b_{4}(M)\leq b_{6}(M)$, as required. ∎
Next, we prove two inequalities on the Betti numbers of a closed symplectic manifold admitting a Hamiltonian $S^{1}$-action, under the assumption that there is no fixed components of positive dimension divisible by $4$. Under the assumption of isolated points, the inequalities were proved by Cho [3, Theorem 1.2]. Both inequalities may be considered as evidence for the unimodality of even Betti numbers of manifolds admitting a Hamiltonian $S^{1}$-action, in the sense that they would be implied by the unimodality. In particular, they imply that the possibilities for any particular Betti number of the form $b_{k}$ with $k=2n\mod 4$ are bounded, if all of the other Betti numbers are fixed, provided $k\neq n$.
Proposition 5.4.
Suppose that $(M,\omega)$ is a closed symplectic manifold of dimension $2n$ having a Hamiltonian $S^{1}$-action, such that there is no fixed component of positive dimension divisible by $4$. Then,
  1. 1.
    If $\dim(M)=0\mod 8$ then:
    $0<1-b_{2}(M)+b_{4}(M)-\dots+b_{n}(M).$
  2. 2.
    If $\dim(M)=4\mod 8$ then:
    $1-b_{2}(M)+b_{4}(M)-\dots-b_{n}(M)\leq 0$
Proof.
By Proposition 3.2 the signature of $M$ is equal to the alternating sum of its even degree Betti numbers.
The first inequality, follows $-b_{n}(M)<\sigma(M)$, after putting all terms to the left hand side, applying Poincaré duality and dividing by $2$. In this case the inequality is strict because the equality case is equivalent to the symplectic form being negative definite, and the top power of the symplectic form is positive.
The second inequality then follows from the inequality $\sigma(M)\leq b_{n}(M)$, where $n=\frac{\dim(M)}{2}$, after putting all terms to the right hand side, applying Poincaré duality and dividing by $2$.
We finally state a proposition regarding the Betti numbers of spin manifolds admitting a Hamiltonian circle action.
Proposition 5.5.
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$ with a Hamiltonian $S^{1}$-action having no fixed component of positive dimension divisible by $4$. Suppose that $n=2\mod 4$ and $M$ is spin, and $b_{2i}(M)=1$ for $2i\neq n$. Then $b_{n}(M)=2\mod 16$.
Proof.
By Proposition 3.2 the signature of $M$ is equal to the alternating sum of its even degree Betti numbers, by the assumptions this simplifies as $2-b_{n}(M)$. By the Rokhlin-Ochanine theorem [27, 24], this quantity is divisible by $16$, the result follows. ∎
The proposition has the following corollary.
Corollary 5.6.
There is no spin fake projective space of dimension $48k+$ 4 having a Hamiltonian $S^{1}$-action with isolated fixed points.